# LongVILA: Scaling Long-Context Visual Language Models for Long Videos

### v1

@@ Abstract
Long-context capability is critical for multi-modal foundation models. We introduce LongVILA, a full-stack solution for long-context vision-language models, including system, model training, and dataset development. On the system side, we introduce the first Multi-Modal Sequence Parallelism (MM-SP) system that enables long-context training and inference, enabling 2M context length training on 256 GPUs. MM-SP is also efficient, being 2.1×2.1$\times$2.1 × - 5.7×5.7$\times$5.7 × faster than Ring-Style Sequence Parallelism and $\times1$ 1.4×1.4$\times$1.4 × faster than Megatron-LM in text-only settings. Moreover, it seamlessly integrates with Hugging Face Transformers. For model training, we propose a five-stage pipeline comprising alignment, pre-training, context extension, and long-short joint supervised fine-tuning. Regarding datasets, we meticulously construct large-scale visual language pre-training datasets and long video instruction-following datasets to support our multi-stage training process. The full-stack solution extends the feasible frame number of VILA by a factor of 128 (from 8 to 1024 frames) and improves long video captioning score from 2.00 to 3.26 (1.6x), achieving 99.5% accuracy in 1400-frames video (274k context length) “needle in a haystack”. LongVILA-8B also demonstrates a consistent improvement in performance on long videos within the VideoMME benchmark as the video frames increase.
@@ 1 Introduction
The importance of Multi-Modal and Long Context From using customized machine learning models to Artificial General Intelligence (AGI), AI research is moving towards a unified solution for everything. To this end, the most exciting breakthrough of foundation model research is unlocking new abilities (Wei et al., 2022), which indicates that the model offers new potential to solve both existing challenges and unforeseen applications. Two important directions of unlocking abilities are multi-modality and long-context.
@@
First, a foundation model supporting more modalities can take more flexible input signals so that people can interact with the model in more diverse manners, e.g., GPT-4o-like multi-modal chatbot, multi-modal web agent (Koh et al., 2024), and real-world robotics foundation model (Brohan et al., 2022; 2023; Padalkar et al., 2023). The longer context enables models to process more information, e.g., long documents, repo-level codebase, and hour-length video, which similarly provides required features to more real-world applications. Since we expect the foundation model to work as a general solution for everything, supporting these two features is highly desirable to achieve real AGI.
Recently, there are many works that enable VLM for short context(Lin et al., 2023b), as well as LLM for long context(Chen et al., 2023; Li et al., 2023a). However, there aren’t much work focusing on Multi-Modal + Long Context at the same time. LWM(Liu et al., 2024a) proposes Ring Attention to extend context length, however the discrete tokens in vision part lead to low performance of image/video understanding although it simplifies the system design. LongVA(Zhang et al., 2024a) proposes training long context LLM and vision encoder on short videos and extend context window during inference. It can help increase context window for VLM but hits limit.
The Necessity of Full-Stack Design in Long Multi-Modal Language Model Compared to the traditional machine learning algorithms (Jordan & Mitchell, 2015) Large Language Model (LLM) training is usually a more systematic project requiring more data engineering (Betker et al., 2023; Ouyang et al., 2022; Zhou et al., 2024) and hardware-software co-design (Lepikhin et al., 2020; Chowdhery et al., 2023; Shoeybi et al., 2019; Brown et al., 2020; Dehghani et al., 2023). Moreover, compared with text-only LLM, the multi-modal foundation model and long-context model both introduce more challenges requiring the multi-component co-design: (1) multi-modal foundation models (e.g., LLaVA (Liu et al., 2023b), and Flamingo (Alayrac et al., 2022)) usually have different model architectures with the text-only LLMs, so that we might need a different distributed training strategy to support the large multi-modal transformer architecture; (2) long context modeling requires both long context data to unlock the ability (Fu et al., 2024b; Chen et al., 2023) and also the infrastructure to make memory-intense long context training possible (Li et al., 2021; Jacobs et al., 2023; Li et al., 2023b). As a result, if we want to introduce both long-context and multi-modal abilities into the foundation model, a careful full-stack design is required to scale the model size and context length up together.
@@
Figure 1: The training pipeline for the LongVILA model extends existing VLM models to support long video processing. Using VILA-1.5 (Lin et al., 2023b) as an example, the process begins with alignment, pre-training, and supervised fine-tuning, in Stages 1 through 3. In Stage 4, the model undergoes mid-training for context extension using the VILA model. Finally, in Stage 5, the model is fine-tuned for long video understanding with MM-specific supervised pretraining (MM-SP), resulting in the LongVILA model. While this example uses VILA, the process is adaptable to other VLMs. Note that we train LongVILA models for up to 1024 video frames in Stage 5.
@@
LongVILA: All-in-One Long-Context Visual Language Model Scaling In this work, we devote ourselves to an all-in-one solution for the training and inference of long-context visual language model. In training, we need to collect (1) a large amount of high-quality visual language data to conduct large-scale pre-training; and (2) high-quality long-context visual language instruction following data to unlock the long-range modeling ability. In addition to the data, we must implement the efficient and easy-to-use infrastructure, i.e., Multi-Modal Sequence Parallelism (MM-SP), supporting (1) multi-modal LLM architecture and (2) memory-intense long context transformer training. From the modeling perspective, we conduct a five-stage training curriculum: (1) multi-modal alignment, (2) large-scale pre-training, (3) short supervised fine-tuning, (4) context extension for LLMs, and (5) long supervised fine-tuning. For inference, the memory usage of KV cache will also be a bottleneck when the sequence length is very long, we thus implement the inference mode of our MM-SP to support long context multi-modal language deployment.
Contributions The contributions of our work can be summarized as follows:
  • From the system level, we proposed Multi-Modal Sequence Parallelism(MM-SP), the first open-sourced sequence parallelism implementation supporting the training and inference of long-context multi-modal language models. Our system is efficient and extremely easy to use (i.e., natively compatible with Huggingface Transformers [^1] https://huggingface.co/docs/transformers/en/index).
  • From the data level, we took both large-scale visual pre-training and long-context multi-modal scaling into consideration. Our data receipt enables our model to generalize well on various domains with 256k tokens.
  • We developed a working training recipe to turn a short context LLM and vision encoder into a long context VLM.
  • As a result, our model achieved competitive results on both video caption and video QA tasks at varied context lengths, which outperforms state-of-the-art open-sourced models by a large margin.
@@ 2 Background and Related work
Figure 2: Scaling the context length can improve multi-modality accuracy, including both Video-QA on VideoMME and Video Captioning on LongVILA-Caption.
Visual language model architecture For VLMs, there are two mainstream designs, i.e., encoder-decoder VLM (e.g., LLaVA (Liu et al., 2023b), PaLM-E (Driess et al., 2023)) and decoder-only VLM (e.g., Fuyu (Bavishi et al., 2023), Chameleon (Team, 2024)).
  • Decoder-only VLMs still lag behind the encoder-decoder VLMs although there are some positive signals showing the potential of decoder-only VLMs. There might be various reasons behind this results.
@@
  • Encoder-Decoder VLMs connect the vision encoder and LLM decoder via the multi-modal projector. Some of the multi-modal projectors (e.g., spatial pooling, Q-former) can significantly reduce the number of tokens for each image or video frame, so that greatly reduce the computational cost of the LLM decoder. However, decoder-only LLMs take raw patches as input without hierarchical token pooling by default, so that make it more challenging to reduce the number of tokens for each image or frame. In this work, we follow VILA (Lin et al., 2023b) as the starting point. Note that there are some other enhanced variants of VILA, including VILA2 (Fang et al., 2024) for better performance and the X-VILA (Ye et al., 2024) for cross-modality understanding, reasoning, and generation. We follow the plain VILA-1.5 version for model architecture and training pipeline.
@@
Sequence parallelism and hybrid strategy Long-context training examples typically produce activation beyond the available memory of one device. To tackle this memory issue, multiple systems have employed the sequence parallelism paradigm in text-only LLMs community, which distributes a single sequence into multiple devices. In particular, Ring-style systems Li et al. (2021; 2023b); Liu et al. (2023a) uses Point-to-Point (P2P) communication primitives to collectively compute the attention module, while DeepSpeed-Ulysses Jacobs et al. (2023) uses All-to-All (A2A) primitive to change the sharding dimension between sequence dimension and attention head dimension to compute the attention module. Our paper is the first to design and implement a sequence parallelism system for visual language models. Furthermore, we carefully study the A2A and P2P primitives and design a novel hybrid sequence parallelism strategy to optimize training throughput. (Gu et al., 2024) is a concurrent work which also investigates the integration of these two SP strategies. However, it is limited to vanilla LLM training and does not extend to multi-modal scenarios due to the lack of support for complex attention mask inputs and variable-length input sequences.
3 System
@@
Long-context VLMs training induces excessive memory footprints. For instance, during our Stage 4 long video training (Figure  1), a single sequence consists of 128K tokens, whose activation consume more memory than available in a single GPU. The most popular open-sourced solution, fully sharded data parallelism does not distribute the activation produced by a single sequence, and thus can not support our use case. Thus, we develop our own system based on sequence parallelism (Li et al., 2021; 2023b; Liu et al., 2023a; Jacobs et al., 2023), commonly implemented in existing foundation model system to optimize text-only LLM training. However, we find that existing systems are not efficient and not scalable enough to support our long-context VLMs workload.
@@
Figure 3: MM-SP comprises of our novel 2D attention technique that uses intra-node All2All and inter-node P2P to communicate keys and values (KV), and a novel sharding strategy that balances the computation of both the image encoder and the language modeling stage. In contrast, Ring-style SP (Liu et al., 2023a; Li et al., 2021; 2023b) employs P2P in both an intra-node and an inter-node setting, causing excessive communication overhead. In addition, they are developed in text-only modality, where the workload of an image encoder is not optimized. Intra-node and inter-node bandwidths are measured in internal H100 clusters.
@@ 3.1 Limitations of Existing Systems
In this section, we discuss the inefficiency of current sequence-parallel systems when applying multi-modal LLMs.
Modality heterogeneity In text-only LLMs, sequences are processed by a single tokenizer into tokens, where it is straightforward to distribute tokens evenly into multiple GPUs. However, VLMs include an encoder architecture, where a non-text data is usually first represented by a single place-holder token (e.g. $<$img$>$), and encoded with various real tokens during training. For example, a video frame usually take around 256 tokens (Lin et al., 2023b). Due to the different processing in visual and text modality, a simple implementation that treats place-holder tokens equally as a text token results in a workload imbalance between GPUs (Figure  3).
Network heterogeneity Our multi-modality comprises extremely long videos (Figure 1), which requires employing sequence parallelism in a multi-node setting. In a multi-node setting, inter-node and intra-node network bandwidth differs significantly. For example, the NVIDIA DGX H100 utilizes NVLink at 900 GB/s for intra-node GPU communication and InfiniBand at 50 GB/s for inter-node GPU communication (single path), resulting in an 18$\times$ difference in bandwidth. Previous work, Ring-Style sequence parallelism (Li et al., 2021; 2023b; Liu et al., 2023a; Zhu, 2023) ignores the heterogeneous networking feature on GPUs and utilizes P2P communication in both inter-node and intra-node settings. This design induces excessive communication costs where they usually attempt to overlap them into computation. However, we found that this design cannot always hide the overhead, and even slows down the computation kernel (Table 1).
Table 1: The forward and backward attention kernel wall-clock time with or without the overlapping design (Unit: $\mu s$). The communication overlap designs in Ring-style SP slows down the attention kernel by acquiring streaming multiprocessor (SM).
@@
Limited maximal sequence length supported DeepSpeed-Ulysses (Jacobs et al., 2023) offers a potential solution to address the communication problem in ring-style SP. It uses All-to-All communication primitives, which induces less communication volumn. Unfortunately, this use of All-to-All primitives come with costs - its design is bind to paralleling in the attention head dimension, instead of the sequence dimension in the attention computation. Thus, simply using DeepSpeed-Ulysses cannot scale beyond the number of attention heads. For example, Llama-3 8B model employs GQA with 8 Key-Value heads, limiting the maximum sequence parallelism degree to 8. Even further employing the replication approach for Key-Value heads with additional communication overhead (Li et al., 2023b), the highest achievable sequence parallelism degree is still restricted to 32 (number of Query heads). This is not sufficient when handling extremely long sequences (e.g, full movies).
@@ 3.2 Multi-Modal Sequence Parallelism Training Mode
Observing problems in existing systems, we summarize that an ideal multi-modal sequence parallelism should achieve efficiency and scalability, by addressing the modality heterogeneity and the network heterogeneity, and also be able to scale beyond the number of attention heads.
@@
Figure 4: Workflow of Multi-Modal Sequence Parallelism, where Batch Size (bs) = 4 and Sequence Parallel Size (SP_Size) = 3. To enable sequence parallelism for multi-modal inputs, we implement a tailored sharding strategy that achieves balanced workload distribution and is compatible with our 2D-attention mechanism.
@@
MM-SP workflow To address the issue of modality heterogeneity, we propose a two-stage sharding strategy that optimizes the compute workload for both the image encoding and language modeling stages. As depicted in Figure 4, the process begins with evenly distributing images (e.g., video frames) across devices within the SP process group, thereby achieving load balancing during the image encoding stage. Subsequently, we gather the global vision and text inputs for token-level sharding in the second stage. To support ring-based attention, we extend sequences with arbitrary dummy tokens, ensuring each sequence can be divided by the ring-based SP degree. This modification maintains consistency with the original version by adjusting label inputs to ignore the padded tokens when calculating loss. We employ a balanced sharding strategy to distribute the context to each rank from both ends, ensuring equal computation across ranks. The advantage of this strategy will be demonstrated later (Table 4). Since this redistribution is performed only once during training, the overhead is minimal. Finally, the balanced local inputs are fed into the LLM backbone, utilizing the 2D-attention mechanism to achieve efficient sequence parallelism.
@@
Figure 5: Comparison of RingAttention, ZigZag-RingAttn, and 2D-Attention. The blue block represents there is commutation between QKV and the black frame indicates local attention computation on each SP group rank (SP_Size=3).
@@
2D-Attention mechanism To address the network heterogeneity and achieve scalability, we leverage the advantages in ring-style SP and Ulysses SP. More specifically, we treat paralleling across the sequence dimension or the attention head dimension as a “1D SP”. Our method parallelizes computation across both attention head and sequence dimension. In particular, we convert 1D SP process group into a 2D mesh, including independent Ring (P2P) process groups and Ulysses (A2A) process groups. For example, as shown in Figure 3 left, to enable an 8-degree sequence parallelism across 2 nodes, we can construct a 4$\times$2 communication mesh using 2D-SP. In this configuration, the A2A process group, with a size of 4, distributes the QKV tensors according to the head dimension and re-partitions them along the sequence dimension within each node. Meanwhile, the P2P process group, with a size of 2, transfers the partitioned KV chunks between nodes. Besides, to further illustrate how ZigZag-RingAttn balances computation and how 2D-attention works, we depict the attention computation schedule using different methods in Figure 5. In this example, the sequence length is 8 and the global SP degree is 4. Due to the triangular structure of causal attention computations, RingAttention experiences a computation imbalance, with rank 0 becoming idle after the first round while rank 3 continues computing through all stages. ZigZag-RingAttn  addresses this by reordering input sequence tokens along the sequence dimension to achieve load balance. Our 2D-attention mechanism uses a ring parallel degree of 2 and a head parallel degree of 2, resulting in an equivalent global sequence parallel degree of 4. This approach also incorporates a workload balance strategy within the ring-based process group. Additionally, it employs the All2All operation to distribute the QKV tensors based on the head dimension across devices, ensuring efficient and balanced computation.

### v2

@@ Abstract
Long-context capability is critical for multi-modal foundation models. We introduce LongVILA, a full-stack solution for long-context vision-language models, including system, model training, and dataset development. On the system side, we introduce the first long-context Multi-Modal Sequence Parallelism (MM-SP) system that enables long training and inference, enabling 2M context length training on 256 GPUs without any gradient checkpointing. MM-SP is 2.1×2.1$\times$2.1 × - 5.7×5.7$\times$5.7 × faster than ring sequence parallelism and $\times1$ 1.4×1.4$\times$1.4 × faster than Megatron context parallelism + tensor parallelism in text-only settings. Moreover, it seamlessly integrates with Hugging Face Transformers. For model training, we propose a five-stage pipeline comprising alignment, pre-training, short supervised fine-tuning, context extension, and long supervised fine-tuning. On datasets, we construct large-scale visual language pre-training datasets and long video instruction-following datasets to support our multi-stage training process. LongVILA extends the number of frames of VILA from 8 to 1024, and improves the long video captioning score from 2.00 to 3.26 (1.6x), achieving 99.5% accuracy in 1400-frames video (274k context length) needle-in-a-haystack. LongVILA-8B demonstrates consistent accuracy improvements on long videos in the VideoMME benchmark as the number of frames increases.
@@ 1 Introduction
Integrating multi-modal understanding with long-context capability is important. A foundation model supporting more modalities can take more flexible input signals so that people can interact with the model in more diverse manners, e.g., GPT-4o-like multi-modal chatbot, multi-modal web agent (Koh et al., 2024), and real-world robotics foundation model (Brohan et al., 2022; 2023; Padalkar et al., 2023). Longer context enables models to process more information, e.g., long documents, repo-level codebase, and hour-length video, which similarly provides required features to more real-world applications.
While some works have enabled long-context Vision-Language Models (VLMs) (Lin et al., 2023b; Weng et al., 2024), they employ simplified approaches rather than offering a comprehensive solution. For instance, LongVA (Zhang et al., 2024a) depends on long-context LLMs and trains models on short-context data. LongVLM (Weng et al., 2024) utilizes token compression to circumvent context extension. These approaches sidestep more challenging issues, such as the development of a robust long-context multi-modal training framework and corresponding dataset design.
A full-stack design is crucial for long-context Vision-Language Models (VLMs). Training large models is typically a complex, systematic endeavor that demands both data engineering (Betker et al., 2023; Ouyang et al., 2022; Zhou et al., 2024) and system-software co-design (Lepikhin et al., 2020; Chowdhery et al., 2023; Shoeybi et al., 2019; Brown et al., 2020; Dehghani et al., 2023). Unlike text-only LLMs, VLMs (e.g., LLaVA (Liu et al., 2023b)) often require distinct model architectures and flexible distributed training strategies. Additionally, long-context modeling necessitates not only long-context data to fully utilize the model’s capabilities (Fu et al., 2024b; Chen et al., 2023) but also infrastructure capable of supporting memory-intensive long-context training (Li et al., 2021; Jacobs et al., 2023; Li et al., 2023a). Therefore, a meticulously planned full-stack design, encompassing system, data, and pipeline, is indispensable for long-context VLMs.
@@
Figure 1: The LongVILA training pipeline for long video understanding. In Stages 1 through 3, the process starts with alignment, pre-training, and supervised fine-tuning. In Stage 4, the model undergoes mid-training context extension. Finally, in Stage 5, the model is fine-tuned for long video understanding with Multi-Modal Sequence Parallelism (MM-SP). While this example uses VILA, the process is adaptable to other VLMs. We train LongVILA models to support 1024 video frames in Stage 5.
@@
In this work, we introduce LongVILA, a comprehensive solution for long-context VLMs. For training infrastructure, we establish an efficient and user-friendly framework, namely Multi-Modal Sequence Parallelism (MM-SP), which supports training memory-intensive long-context VLMs. For training pipeline, we implement a five-stage training curriculum as Figure 1: (1) multi-modal alignment, (2) large-scale pre-training, (3) short supervised fine-tuning, (4) context extension for LLMs, and (5) long supervised fine-tuning. For inference, MM-SP addresses the challenge of KV cache memory usage, which becomes a bottleneck with very long sequences, by deploying an inference mode within our MM-SP framework to support long-context multi-modal language deployment. For training data, we focus on two key requirements: (1) the collection of a large volume of high-quality visual-language data for large-scale pre-training, and (2) the acquisition of high-quality long-context visual-language instruction-following data to unlock long-range modeling capabilities.
By efficiently increasing the number of video frames using LongVILA, the performance on VideoMME (Fu et al., 2024a) and long video captioning tasks consistently improves (Figure 2). The LongVILA model, trained on 1024 frames, achieves 99.5% accuracy in the needle-in-a-haystack experiments with 1400 frames, equivalent to a context length of 274k tokens. Our MM-SP system can efficiently scale the context length up to 2 million tokens without gradient checkpointing, achieving 2.1$\times$ to 5.7$\times$ speedup compared to ring sequence parallelism, and 1.1$\times$ to 1.4$\times$ compared to Megatron context parallelism + tensor parallelism.
@@ 2 Background and Related work
Visual language model architecture.
@@
There are two predominant designs for VLMs: the encoder-decoder architecture (e.g.,, LLaVA (Liu et al., 2023b), PaLM-E (Driess et al., 2023)) and the decoder-only architecture (e.g.,, Fuyu (Bavishi et al., 2023), Chameleon (Team, 2024)). Encoder-Decoder VLMs connect the vision encoder to the LLM decoder through a multi-modal projector. Certain multi-modal projectors, such as spatial pooling and Q-former, significantly reduce the number of tokens per image or video frame, thereby lowering the computational burden on the LLM decoder. In contrast, decoder-only LLMs typically process raw patches as input without hierarchical token pooling, making it more challenging to reduce the token count for each image or frame. In this work, we build on VILA (Lin et al., 2023b) as our foundation. It is worth noting that enhanced variants of VILA exist, such as VILA2(Fang et al., 2024) for improved performance and X-VILA(Ye et al., 2024) for cross-modality understanding, reasoning, and generation. For our model architecture and training pipeline, we adhere to the standard VILA-1.5 version.
Figure 2: Scaling the context length can enhance multi-modality accuracy, improving performance on both Video-QA tasks with VideoMME and video captioning with LongVILA-Caption.
Sequence parallelism and hybrid strategy.
@@
Long-context training examples often exceed the memory capacity of a single device. To address this issue, the sequence parallelism paradigm has been widely adopted in the text-only LLM community, distributing a single sequence across multiple devices. Specifically, Ring-style systems Li et al. (2021; 2023a); Liu et al. (2023a) use Point-to-Point (P2P) communication primitives to collectively compute the attention module, while DeepSpeed-Ulysses Jacobs et al. (2023) employs an All-to-All (A2A) primitive to alternate between sharding the sequence dimension and the attention head dimension during attention computation. Our work is the first to design and implement a sequence parallelism system for visual language models. Moreover, we have thoroughly examined both A2A and P2P primitives and developed a novel hybrid sequence parallelism strategy to optimize training throughput. Concurrent work, such as (Gu et al., 2024), also explores integrating these two SP strategies. However, it is limited to conventional LLM training and does not extend to multi-modal scenarios due to the lack of support for complex attention masks and variable-length input sequences.
3 Multi-Modal Sequence Parallelism
@@
Training long-context Vision-Language Models (VLMs) results in substantial memory demands. For example, in our Stage 5 long video training (Figure 1), a single sequence comprises around 200K tokens for 1024 video frames, which exceeds the memory capacity of a single GPU. The most widely used open-source solution, fully sharded data parallelism, does not distribute the activations generated by a single sequence, making it unsuitable for our needs. Consequently, we developed a custom system based on sequence parallelism(Li et al., 2021; 2023a; Liu et al., 2023a; Jacobs et al., 2023), a technique commonly employed in existing foundation model systems to optimize text-only LLM training. However, we discovered that existing systems are neither efficient nor scalable enough to handle our long-context VLM workloads.
@@
Figure 3: MM-SP incorporates our innovative 2D attention technique, which utilizes intra-node All-to-All (All2All) and inter-node Point-to-Point (P2P) communication to transfer keys and values (KV). It also includes a novel sharding strategy that balances the computational load between the image encoder and the language modeling stages. In contrast, Ring-style SP (Liu et al., 2023a; Li et al., 2021; 2023a) relies on P2P communication for both intra-node and inter-node settings, leading to significant communication overhead. Additionally, these systems are designed for text-only modalities, without optimization for the workload of an image encoder. Intra-node and inter-node bandwidths were measured in internal H100 clusters.
@@ 3.1 Limitations of Existing Systems
In this section, we address the inefficiencies of current sequence-parallel systems when applied to multi-modal LLMs.
Modality heterogeneity. In text-only LLMs, sequences are processed by a single tokenizer into tokens, allowing for straightforward distribution of tokens across multiple GPUs. However, VLMs incorporate an encoder architecture where non-text data is initially represented by a placeholder token (e.g., $<$img$>$) and subsequently encoded into multiple real tokens during training. For instance, a single video frame typically requires around 256 tokens (Lin et al., 2023b). Due to the differing processing requirements of visual and text modalities, a simplistic implementation that treats placeholder tokens the same as text tokens leads to an imbalance in GPU workloads (Figure 3).
Networking heterogeneity. Our multi-modality comprises extremely long videos (Figure 1), which requires employing sequence parallelism in a multi-node setting. In a multi-node setting, inter-node and intra-node network bandwidth differs significantly. For example, the NVIDIA DGX H100 utilizes NVLink at 900 GB/s for intra-node GPU communication and InfiniBand at 50 GB/s for inter-node GPU communication (single path), resulting in an 18$\times$ difference in bandwidth. Previous work, Ring-Style sequence parallelism (Li et al., 2021; 2023a; Liu et al., 2023a; Zhu, 2023) ignores the heterogeneous networking feature on GPUs and utilizes P2P communication in both inter-node and intra-node settings. This design induces excessive communication costs where they usually attempt to overlap them into computation. However, we found that this design cannot always hide the overhead, and even slows down the computation kernel (Table 1).
Our multi-modality approach involves processing extremely long videos, necessitating the use of sequence parallelism in a multi-node setting. In such environments, there is a significant disparity between inter-node and intra-node network bandwidth. For instance, the NVIDIA DGX H100 employs NVLink at 900 GB/s for intra-node GPU communication and InfiniBand at 50 GB/s for inter-node GPU communication (single path), resulting in an 18-fold difference in bandwidth. Previous work on Ring-Style sequence parallelism (Li et al., 2021; 2023a; Liu et al., 2023a; Zhu, 2023) overlooks the heterogeneous nature of GPU networking and relies on P2P communication for both inter-node and intra-node settings. This approach incurs substantial communication costs, which are typically overlapped with computation. However, we found that this design does not consistently mitigate the overhead and can even slow down the computation kernel (Table 1).
Table 1: The forward and backward attention kernel wall-clock time with or without the overlapping design (Unit: $\mu$s). The communication overlap design in Ring-style SP slows down the attention kernel by occupying streaming multiprocessor (SM) resources.
@@
Limited maximal sequence length. DeepSpeed-Ulysses (Jacobs et al., 2023) presents a potential solution to the communication challenges in ring-style sequence parallelism by employing All-to-All communication primitives, which reduce the overall communication volume. However, this approach has its limitations. The design relies on parallelizing along the attention head dimension rather than the sequence dimension during attention computation. As a result, DeepSpeed-Ulysses cannot scale effectively beyond the number of attention heads. For instance, the Llama-3 8B model uses Grouped Query Attention (GQA) with 8 Key-Value heads, which restricts the maximum sequence parallelism degree to 8. Even when using replication for Key-Value heads, which introduces additional communication overhead (Li et al., 2023a), the highest achievable sequence parallelism degree is still limited to 32 (the number of Query heads). This constraint is insufficient for handling extremely long sequences, such as full-length movies.
@@ 3.2 Multi-Modal Sequence Parallelism Training Mode
After identifying the limitations in existing systems, we conclude that an ideal multi-modal sequence parallelism approach should prioritize efficiency and scalability by addressing both modality and network heterogeneity, and should also be capable of scaling beyond the number of attention heads.
@@
Figure 4: Workflow of Multi-Modal Sequence Parallelism with a Batch Size (bs) of 4 and a Sequence Parallel Size (SP_Size) of 3. To facilitate sequence parallelism for multi-modal inputs, we developed a customized sharding strategy that ensures balanced workload distribution and compatibility with our 2D-attention mechanism.
@@
MM-SP workflow. To address the challenge of modality heterogeneity, we propose a two-stage sharding strategy that optimizes the compute workload for both image encoding and language modeling stages. As illustrated in Figure 4, the process begins by evenly distributing images (e.g., video frames) across devices within the sequence parallelism (SP) process group, thereby achieving load balancing during the image encoding stage. In the second stage, we aggregate global vision and text inputs for token-level sharding. To support ring-based attention, sequences are extended with arbitrary dummy tokens, ensuring that each sequence can be evenly divided according to the ring-based SP degree. This adjustment maintains consistency with the original approach by modifying label inputs to ignore padded tokens during loss calculation. We implement a balanced sharding strategy that distributes the context to each rank from both ends, ensuring equal computation across ranks. The effectiveness of this strategy will be demonstrated later (Table 4). Since this redistribution is performed only once during training, the overhead is minimal. Finally, the balanced local inputs are processed by the LLM backbone, utilizing 2D-attention to achieve efficient sequence parallelism.
@@
Figure 5: Comparison of RingAttention, ZigZag-RingAttn, and 2D-Attention. The blue block indicates communication between QKV, while the black frame represents local attention computation within each SP group rank (SP_Size = 3).
2D attention parallelism. To address network heterogeneity and achieve scalability, we combine the strengths of ring-style sequence parallelism (SP) and Ulysses SP. Specifically, we consider parallelization across the sequence dimension or the attention head dimension as ”1D SP.” Our method extends this by parallelizing computation across both the attention head and sequence dimensions, transforming a 1D SP process group into a 2D mesh comprising independent Ring (P2P) and Ulysses (A2A) process groups. For instance, as illustrated on the left in Figure 3, to enable an 8-degree sequence parallelism across 2 nodes, we construct a 4$\times$2 communication mesh using 2D-SP. In this setup, the A2A process group, with a size of 4, distributes the QKV tensors according to the head dimension and re-partitions them along the sequence dimension within each node. Simultaneously, the P2P process group, with a size of 2, transfers the partitioned KV chunks between nodes.
@@
Additionally, to further explain how ZigZag-RingAttnbalances computation and how the 2D-attention mechanism operates, we depict the attention computation schedule using different methods in Figure 5. In this example, the sequence length is 8 and the global SP degree is 4. Due to the triangular structure of causal attention computations, RingAttention experiences a computation imbalance, where rank 0 becomes idle after the first round while rank 3 continues computing through all stages. ZigZag-RingAttnaddresses this by reordering input sequence tokens along the sequence dimension to achieve load balance. Our 2D-attention mechanism uses a ring parallel degree of 2 and a head parallel degree of 2, resulting in an effective global sequence parallel degree of 4. This approach also incorporates a workload balancing strategy within the ring-based process group and uses the All-to-All operation to distribute QKV tensors across devices based on the head dimension, ensuring efficient and balanced computation.